OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (anonymous):

\[turn~ 5x-3y=-6 ~to~ slope~ intercept~ form~ (y=mx+b)\]\[5x-3y=-6~~\rightarrow~~-3y=-5x-6\rightarrow~~y={3\over5}x-6\]

OpenStudy (anonymous):

\[m=~slope\]\[b=~y-intercept\]

OpenStudy (anonymous):

do you get it? :)

OpenStudy (anonymous):

@chariety101 ?? :)

OpenStudy (anonymous):

yes i get it so far

OpenStudy (anonymous):

so what is in place of m and b in y=3/5x+2 ??:)

OpenStudy (anonymous):

im correcting that \[−3y=−5x−6~~→~ y={3\over5}x+2\]

OpenStudy (anonymous):

m = 3/5 and b = 2 ?

OpenStudy (anonymous):

yup! :D

OpenStudy (anonymous):

ohhh i get this!

OpenStudy (anonymous):

yay! ^_^

OpenStudy (anonymous):

thanks so much!!

OpenStudy (anonymous):

your welcome! :)...lemme know if you have anymore questions :D

OpenStudy (anonymous):

alright thanks so much(:

OpenStudy (anonymous):

anytime! :)